OpenStudy (anonymous):

Use the Laplace transformation to solve the initial value problem:

OpenStudy (anonymous):

\[\huge y''+6y'+18y=0\]

OpenStudy (anonymous):

I was able to plug everything in and reduce it down to\[\large Y(s)=\frac{ 4s+13 }{ s^2+6s+18 }\] (if someone could check this, that'd be awesome) I don't know where to go from here since the denominator is irreducible.

OpenStudy (anonymous):

Oops... initial values..\[y(0)=4, ~~~y'(0)=-11\]

zepdrix (zepdrix):

Hmm I got -4s-13 in the numerator. Sec checking my work...

zepdrix (zepdrix):

Oh oh but I have to add them to the right side, derp. Ok ya looks good so far :)

OpenStudy (johnweldon1993):

Everything looks good here

zepdrix (zepdrix):

So we have something like...\[\large\rm Y(s)=\frac{4s+13}{(s+3)^2+3^2}\]ya?

zepdrix (zepdrix):

I'll bet we can make an (s+3) in the numerator somehow, and then split this up into (s+3)/(stuff) which turns into cosine, ya? And another fraction that we have to figure out.

OpenStudy (anonymous):

Oh you completed the square. Yeah, that makes sense X)

OpenStudy (johnweldon1993):

@zepdrix for the win lol

zepdrix (zepdrix):

4s+13 = 4s+12 +1 = 4(s+3) +1

OpenStudy (anonymous):

Gotcha, okay still following you!

zepdrix (zepdrix):

\[\large\rm Y(s)=\frac{4(s+3)+1}{(s+3)^2+3^2}\]Giving us,\[\large\rm Y(s)=4\frac{(s+3)}{(s+3)^2+3^2}+\frac{1}{(s+3)^2+3^2}\]And that second term, hmmm... Does that look like anything useful? I can't remember these things sometimes. Oh it's close to sine, isn't it?

zepdrix (zepdrix):

I mean, it's close to sine, with an exponential shift or whatever (from the s+3)

OpenStudy (anonymous):

Isn't it close to e^at?

zepdrix (zepdrix):

\[\large\rm Y(s)=4\frac{(s+3)}{(s+3)^2+3^2}+\frac{1}{3^2}\frac{3^2}{(s+3)^2+3^2}\]

OpenStudy (anonymous):

Oh wait, e^at doesn't have a^2

zepdrix (zepdrix):

Or even an s squared right? :D

OpenStudy (anonymous):

True :P Thank you! This helped a ton